# Unraveling-induced entanglement phase transition in diffusive trajectories of continuously monitored noninteracting fermionic systems

### v1

@@ Acknowledgements.
Acknowledgements.— We thank Marcel Cech, Gabriele Perfetto and Chris Nill for fruitful discussions. We acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project No. 435696605 and through the Research Unit FOR 5413/1, Grant No. 465199066, and through the Research Unit FOR 5522/1, Grant No. 499180199. This project has also received funding from the European Union’s Horizon Europe research and innovation program under Grant Agreement No. 101046968 (BRISQ). F.C. is indebted to the Baden-Württemberg Stiftung for the financial support of this research project by the Eliteprogramme for Postdocs.
SUPPLEMENTAL MATERIAL
Unraveling-induced entanglement phase transition in diffusive trajectories of continuously monitored noninteracting fermionic systems
Moritz Eissler,1 Igor Lesanovsky1,2, and Federico Carollo1
1Institut für Theoretische Physik, Universität Tübingen,
Auf der Morgenstelle 14, 72076 Tübingen, Germany
2School of Physics and Astronomy and Centre for the Mathematics and Theoretical Physics of Quantum Non-Equilibrium Systems, The University of Nottingham, Nottingham, NG7 2RD, United Kingdom
I Numerical Implementation
@@
I.1 Representation of the state and covariance matrix
@@
$C^{\dagger}_{k}=\sum_{j=1}^{L}\tilde{U}_{jk}c^{\dagger}_{j},\,\qquad C_{k}=\sum_{j=1}^{L}\tilde{U}^{*}_{jk}c_{j}\,,$ (S1)
@@
$D_{mn}=\langle\psi_{t}|c^{\dagger}_{m}c_{n}|\psi_{t}\rangle=\sum_{j,k=1}^{L}\langle 0|C_{1}\dots C_{N}\tilde{U}^{*}_{mj}\tilde{U}_{nk}C^{\dagger}_{j}C_{k}C^{\dagger}_{N}\dots C^{\dagger}_{1}|0\rangle.$
@@
I.2 Time evolution of the state
@@
$K_{t}=-iH{\rm d}t-\sum_{i=1}^{L}\langle n_{i}\rangle_{t}\left(\sqrt{\gamma_{\mathrm{I}}}{\rm d}W_{t}^{i}+\gamma_{\mathrm{I}}\langle n_{i}\rangle_{t}{\rm d}t\right)+\sum_{i=1}^{L}n_{i}\left(\gamma_{\mathrm{I}}\left(2\langle n_{i}\rangle_{t}-1\right){\rm d}t-i\sqrt{\gamma_{\mathrm{II}}}{\rm d}V_{t}^{i}+\sqrt{\gamma_{\mathrm{I}}}{\rm d}W_{t}^{i}\right)\,.$ (S2)
In the numerical implementation of the above equation, we consider a small, but finite, ${\rm d}t$ and we generate Wiener processes by considering ${\rm d}W_{t}^{i}$ and ${\rm d}V_{t}^{i}$ to be random Gaussian numbers with average zero and variance $\sqrt{{\rm d}t}$. For our numerical simulations, we use $\gamma{\rm d}t=0.1$.
@@
$e^{K_{t}}\ket{\psi_{t}}=e^{K_{t}}\prod_{k=1}^{N}\left(\sum_{j=1}^{L}U_{jk}(t)c^{\dagger}_{j}\right)|0\rangle=\prod_{k=1}^{N}\left(\sum_{j=1}^{L}U_{jk}(t)e^{K_{t}}c^{\dagger}_{j}e^{-K_{t}}\right)|0\rangle\,.$
@@
$U_{jk}(t+{\rm d}t)=\left[Me^{-ih{\rm d}t}U(t)\right]_{jk}\,,\quad\mbox{with}\quad M_{jk}=\delta_{jk}e^{\left[\gamma_{\mathrm{I}}\left(2\langle n_{k}\rangle_{t}-1\right){\rm d}t-i\sqrt{\gamma_{\mathrm{II}}}{\rm d}V_{t}^{k}+\sqrt{\gamma_{\mathrm{I}}}{\rm d}W_{t}^{k}\right]}\,.$
@@
$S=-\sum_{j=1}^{\ell}\left[\lambda_{j}^{(A)}\log\left(\lambda_{j}^{(A)}\right)+\left(1-\lambda_{j}^{(A)}\right)\log\left(1-\lambda_{j}^{(A)}\right)\right].$ (S3)

### v2

@@ Acknowledgements.
We thank Marcel Cech, Gabriele Perfetto and Chris Nill for fruitful discussions. We are further grateful to Xhek Turkeshi for useful discussions and for drawing our attention to recent literature. We acknowledge funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Project No. 435696605 and through the Research Unit FOR 5413/1, Grant No. 465199066, through the Research Unit FOR 5522/1, Grant No. 499180199 and from the state of Baden-Württemberg through bwHPC grant no. INST 40/575-1 FUGG (JUSTUS 2 cluster). This project has also received funding from the European Union’s Horizon Europe research and innovation program under Grant Agreement No. 101046968 (BRISQ). F.C. is indebted to the Baden-Württemberg Stiftung for the financial support of this research project by the Eliteprogramme for Postdocs. This work was supported by the QuantERA II programme (project CoQuaDis, DFG Grant No. 532763411) that has received funding from the EU H2020 research and innovation programme under GA No. 101017733.
Appendix A Numerical implementation
@@
A.1 Representation of the state and covariance matrix
@@
$C^{\dagger}_{k}=\sum_{j=1}^{L}\tilde{U}_{jk}c^{\dagger}_{j},\,\qquad C_{k}=\sum_{j=1}^{L}\tilde{U}^{*}_{jk}c_{j}\,,$ (A1)
@@
$D_{mn}$ $=\langle\psi_{t}|c^{\dagger}_{m}c_{n}|\psi_{t}\rangle$
$=\sum_{j,k=1}^{L}\langle 0|C_{1}\dots C_{N}\tilde{U}^{*}_{mj}\tilde{U}_{nk}C^{\dagger}_{j}C_{k}C^{\dagger}_{N}\dots C^{\dagger}_{1}|0\rangle.$
@@
A.2 Time evolution of the state
@@
$K_{t}=$ $-iH{\rm d}t-\sum_{i=1}^{L}\langle n_{i}\rangle_{t}\left(\sqrt{\gamma_{\mathrm{I}}}{\rm d}W_{t}^{i}+\gamma_{\mathrm{I}}\langle n_{i}\rangle_{t}{\rm d}t\right)$
$+\sum_{i=1}^{L}n_{i}\left(\gamma_{\mathrm{I}}\left(2\langle n_{i}\rangle_{t}-1\right){\rm d}t-i\sqrt{\gamma_{\mathrm{II}}}{\rm d}V_{t}^{i}+\sqrt{\gamma_{\mathrm{I}}}{\rm d}W_{t}^{i}\right)\,.$ (A2)
In the numerical implementation of the above equation, we consider a small, but finite, ${\rm d}t$ and we generate Wiener processes by considering ${\rm d}W_{t}^{i}$ and ${\rm d}V_{t}^{i}$ to be random Gaussian numbers with average zero and variance $\sqrt{{\rm d}t}$. For our numerical simulations, we use $\gamma{\rm d}t=0.1$.It is notable, that while the trotterization (A2) might hold for smaller $\rm{d}t$ it might not work for the one chosen. In order reconsidering this we resimulated the data setting $\rm{d}t=0.05$ and recovered the same expected behaviour for the $\rm{d}t$ chosen as seen in Fig. B3.
@@
$e^{K_{t}}\ket{\psi_{t}}$ $=e^{K_{t}}\prod_{k=1}^{N}\left(\sum_{j=1}^{L}U_{jk}(t)c^{\dagger}_{j}\right)|0\rangle$
$=\prod_{k=1}^{N}\left(\sum_{j=1}^{L}U_{jk}(t)e^{K_{t}}c^{\dagger}_{j}e^{-K_{t}}\right)|0\rangle\,.$
@@
$U_{jk}(t+{\rm d}t)=\left[Me^{-ih{\rm d}t}U(t)\right]_{jk}\,,$
$\mbox{with}\quad M_{jk}=\delta_{jk}e^{\left[\gamma_{\mathrm{I}}\left(2\langle n_{k}\rangle_{t}-1\right){\rm d}t-i\sqrt{\gamma_{\mathrm{II}}}{\rm d}V_{t}^{k}+\sqrt{\gamma_{\mathrm{I}}}{\rm d}W_{t}^{k}\right]}\,.$
@@
$S=-\sum_{j=1}^{\ell}\left[\lambda_{j}^{(A)}\log\left(\lambda_{j}^{(A)}\right)+\left(1-\lambda_{j}^{(A)}\right)\log\left(1-\lambda_{j}^{(A)}\right)\right].$ (A3)
Appendix B Additional results
In this section, we provide additional results, including an explicit fit of the Ansatz to the generated data, additional data for a smaller time-increment, plots regarding both volume-law saturation values and the initial entropy dynamics as well as a more systematic approach to the finite-size scaling.
B.1 Results for a smaller time-step size
In this section we show results for $\gamma\rm{dt}=0.05$ in order to rule out the possibility that the behavior observed in the main text is influenced by a too large Trotterization step. These results are shown in Fig. B1 and display an entanglement entropy behavior compatible with the one discussed in the main text for $\gamma\rm{d}t=0.1$.
Figure B1: Unraveling-induced entanglement phase transition for $\gamma\rm{d}t=0.05.$ (a) Stationary values of the entanglement entropy $\bar{S}_{\infty}(\alpha,L)$ as a function of the logarithm of the system size, for different values of $\alpha$. The inset shows the data in linear scale. For small parameter $\alpha$, $\bar{S}_{\infty}(\alpha,L)$ exhibits an area-law behavior. Upon increasing $\alpha$, $\bar{S}_{\infty}(\alpha,L)$ starts to display a subextensive logarithmic growth with the system size. (b) Estimate of the order parameter $c(\alpha)$ [see Eq. (8)], obtained as $\bar{S}_{\infty}(\alpha,L)/\log L$, for $L\in\left\{100,150,200,250,300,350,400\right\}$. The dashed line represents $c(\alpha)$ from a fit of Eq. (8) using the largest four system sizes considered in panel (a). The inset shows instead the fit of the residual entropy $s_{0}(\alpha)$. We have considered $\lambda=\gamma/2$ and the average is performed over $100$ realizations of the process. (c) Finite-size scaling of the stationary entanglement entropy, assuming BKT universality (see main text and Ref. [27]). The difference $\bar{S}_{\infty}(\alpha,L)-\bar{S}_{\infty}(\alpha_{\rm c},L)$ is plotted against the quantity $(\alpha-\alpha_{\rm c})\log^{2}L$, for $\alpha_{\rm c}=0.8$. The inset shows the bare data for $\bar{S}_{\infty}(\alpha,L)$. We have considered $\lambda=\gamma/2$ and the average is performed over $100$ realizations of the process.
B.2 Volume law, initial entropy dynamics and collapse of stationary entropies
In order to get additional insight onto the entropy dynamics in our model, we also briefly study the transient behavior in the logarithmic phase. While it is known that in the volume law the entanglement entropy grows as $\bar{S}_{t}\propto\sqrt{t}$ [48, 25], we find that in the logarithmic regime it grows as $\propto\log\left(t\right)$, as shown in Fig. B2(a), before saturating to a value $\propto\log\left(L\right)$.
Figure B2: Entanglement entropy dynamics. (a) Entanglement entropy averaged over 100 realizations of the process in the logarithmic phase ($\alpha=0.9$) with a logarithmic x-axis. The plot shows the initial logarithmic-in-time growth of entanglement. The inset displays the bare data for $\alpha=0.9$. (b) Entanglement entropy, $\bar{S}$, averaged over $100$ realizations of the stochastic process in Eq. (6) for $\alpha=1$. Dotted lines represent the entanglement entropy for $\alpha=1$ while the connected lines represent the unitary case $\gamma=0$.
We also look more in detail at the stationary entanglement behavior for two volume-law cases considered: the case of the unravelling with $\alpha=1$ and that of the purely unitary dynamics, or $\gamma=0$ case. We observe that both scenarios yield not only the same volume law character, but the same stationary entanglement as displayed in Fig. B2(b). However, the $\gamma=0$ scenario exhibits an oscillatory behavior around the stationary value while the $\alpha=1$ scenario approaches it directly and remains stable.
Figure B3: Fit of the Ansatz and systematic finite size scaling. (a) Data of the stationary entropy fitted with (8) for all values of $\alpha$ over the logarithm of the system size $L$. The dots represent the numerically obtained data points, and the lines the fit. (b) Total distance $\Delta(y_{i},y_{j})$ of the curves for different critical interpolation values $\alpha_{\rm c}$ extracted assuming BKT universality as exemplary shown in 4. where $\left\{y_{i}\right\}$ are the interpolated y-values $\bar{S}_{\infty}(\alpha,L)-\bar{S}_{\infty}(\alpha_{\rm c},L)$ for each critical parameter $\alpha_{c}$ onto a general grid of x values. The Inset shows the finite size scaling of the stationary entanglement entropy, assuming BKT universality.
Finally, we also provide here a more “objective” approach to the estimate of the critical parameter $\alpha_{\rm c}$, inspired by what has been done in Ref. [38]. To this end, we introduce a function which considers the distances between all pairs of curves obtained by rescaling the numerical data for the different system sizes. Such function solely depends on the free parameter $\alpha_{\rm c}$. More precisely, we consider the normalized distance between curves
$\Delta=\sum_{i,j,i\neq j}\frac{|y_{i}-y_{j}|}{|y_{i}|+|y_{j}|}.$ (B4)
Here, the set $\left\{y_{i}\right\}$ represents the $\ y$-values, defined as $\bar{S}_{\infty}(\alpha,L)-\bar{S}_{\infty}(\alpha_{\rm c},L)$ as in Fig. 4, obtained by interpolation onto a general grid of $x$-values, defined as $\left(\alpha-\alpha_{\mathrm{c}}\right)\log^{2}L$ in order to compare curves obtained for different $L$.
After defining the function, we want to find its minimum upon varying the free parameter $\alpha_{\rm c}$. Such a minimum indicates the value for which the scaling is optimal. According to Fig. B3(b), the best collapse is obtained in the region around $\alpha_{\rm c}\approx 0.875$. Such a value is in line, given the finite-size effects that we have and the interpolation we have to perform due to our finite data set, with the one reported in the main text.